EL019

PI: Lobanov
Title: Lobe-dom Qs

Session: Feb 1997 (session 1)
Frequency: 18 cm
Mode: B
Obs time (UT): EU 16:30(02/03)-04:30(03/03)
Stations: EbJbMcNtOnTrSmShUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Claudio Bortolotti at Mon Mar 3 00:00:24 EST 1997) Success

Onsala (Antonis Polatidis at Mon Mar 3 03:28:40 EST 1997) Minor failures
Due to very strong winds (up to 24 m/s) we could only observe for the first hour of the experiment.

Westerbork (Tony Foley at Tue Mar 4 07:07:12 EST 1997) Success
Looks OK

Shanghai (Huang Xinyong at Wed Mar 5 05:28:09 EST 1997) Success
Clock offset = 0.64 - 0.66 msec. No problems, weather cloudy to fine.

Jodrell (Alastair Gunn at Tue Mar 18 15:20:45 MET 1997) Minor failures
Telescope parked until 18:09 due to high winds, problems with vacuum, data recorded from 18:23 scan. Some problems with vacuum; data missing between 23:00 and end of tape. Very bad interference.

Calibration (Fredrik T Rantakyro at Thu May 8 11:39:08 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name el19.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Walter Alef at Wed Jan 7 12:11:43 MET 1998) Failure
Correlation cancelled due to too many failures.